OpenStudy (anonymous):

How many diagonals are in a polygon with 18 sides

OpenStudy (anonymous):

\[\frac{n(n-3)}{2}\]

OpenStudy (anonymous):

the formula is x(x-3)/2 so 18(18-3)/2 18(15)/2 9(15) 135 diagonals.

OpenStudy (anonymous):

135 diagonals

OpenStudy (anonymous):

\[\left(\begin{matrix}18 \\ 2\end{matrix}\right) - 18\] 135 diagonals.